# Hacking Task Confounder in Meta-Learning

### v1

@@ Abstract
Meta-learning enables rapid generalization to new tasks by learning meta-knowledge from a variety of tasks. It is intuitively assumed that the more tasks a model learns in one training batch, the richer knowledge it acquires, leading to better generalization performance. However, contrary to this intuition, our experiments reveal an unexpected result: adding more tasks within a single batch actually degrades the generalization performance. To explain this unexpected phenomenon, we conduct a Structural Causal Model (SCM) for causal analysis. Our investigation uncovers the presence of spurious correlations between task-specific causal factors and labels in meta-learning. Furthermore, the confounding factors differ across different batches. We refer to these confounding factors as “Task Confounders”. Based on this insight, we propose a plug-and-play Meta-learning Causal Representation Learner (MetaCRL) to eliminate task confounders. It encodes decoupled causal factors from multiple tasks and utilizes an invariant-based bi-level optimization mechanism to ensure their causality for meta-learning. Extensive experiments on various benchmark datasets demonstrate that our work achieves state-of-the-art (SOTA) performance.
Introduction
Meta-learning aims to develop algorithms capable of adapting to previously unseen tasks. To achieve this goal, meta-learning methods are trained on a diverse set of tasks, enabling them to learn meta-knowledge that can be applied to new and related tasks. Moreover, meta-learning has been demonstrated to address numerous inherent challenges in deep learning, such as computational bottlenecks and generalization issues (Du et al. 2020; Li et al. 2018). It is widely used in domains like reinforcement learning (Mitchell et al. 2021), computer vision (Mahadevkar et al. 2022), and robotics (Schrum et al. 2022).
In general, meta-learning can be defined as a bi-level optimization process. In the inner loop, task-specific parameters are independently learned based on meta-parameters. In the outer loop, the meta-parameters are updated by minimizing the average loss across multiple tasks using the learned task-specific parameters. During the meta-training process, each batch’s training set consists of a series of randomly sampled different tasks, with each task containing training samples from various categories. By leveraging shared structures across multiple tasks, the meta-learning model can acquire rich meta-knowledge, leading to great generalization and adaptation (Wang, Zhao, and Li 2021; Song et al. 2022). Therefore, a widely adopted hypothesis is that the more tasks the model learns in a single training batch, the richer knowledge it acquires, and consequently, the better its performance will be (Hospedales et al. 2021; Rivolli et al. 2022). This idea is intuitively reasonable since learning from a broader range of scenarios can help grasp a wealth of knowledge, resembling human cognition.
(a) miniImagenet
(b) Omniglot
(c) tieredImagenet
(d) CIFAR-FS
Figure 1: The empirical results on four benchmark datasets.
However, our experiments yield conflicting results. We sample two sets of non-overlapping tasks, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. Within $\mathcal{T}_{1}$, we divide the tasks into a meta-training set $\mathcal{D}_{1}$ and a meta-testing set $\mathcal{D}_{2}$, while $\mathcal{T}_{2}$ is used solely for separate testing without being split. We train MAML (Finn, Abbeel, and Levine 2017) on $\mathcal{D}_{1}$ with two different batch size settings, i.e., batch size=$B$ and batch size=$2B$. Then we test it on $\mathcal{D}_{2}$ and $\mathcal{T}_{2}$. Intuitively, the model with a larger batch size is expected to perform better. Surprisingly, as shown in Figure 1, when batch size=$2B$, the model exhibits lower accuracy on both $\mathcal{D}_{2}$ and $\mathcal{T}_{2}$ across all four benchmark datasets.
To explore the causes of this phenomenon, we construct a Structural Causal Model (SCM). As shown in Figure 1(b), we denote the distinct causal factors of task $\tau_{i}$ and task $\tau_{j}$ as ${\rm{A}}^{i}$ and ${\rm{A}}^{j}$, as well as the shared causal factors as ${\rm{B}}^{i,j}$. To ensure a generic prior that performs well, meta-learning performs joint learning on all tasks. Thus, the non-overlapping causal factors ${\rm{A}}^{i}$ of $\tau_{i}$ may cause spurious correlations with $\tau_{j}$, and ${\rm{A}}^{j}$ holds the same with $\tau_{i}$. This misleading correlation introduces bias into meta-knowledge, ultimately affecting model generalization. Additionally, it varies across different batches, e.g., the tasks differ in different batches. We identify this confounding factor as “Task Confounder”.
@@ Introduction
(b) meta-learning process in practice
@@
Figure 2: The Structural Causal Model (SCM) regarding two tasks $\tau_{i}$ and $\tau_{j}$, where $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$ are the samples and corresponding labels of these two tasks. The solid line means the causal correlation, and the dotted line means the spurious correlation. (a) is constructed based on the ground-truth causal mechanism, while (b) can be viewed as the inverse process of the generating mechanism.
@@
Inspired by this insight, we propose a plug-and-play meta-learning causal representation learner (MetaCRL) to encode decoupled causal knowledge, thereby eliminating task confounders. It consists of two modules: the disentangling module and the causal module. The former aims to extract causal factors across all tasks and provide a subset of causal factors relevant to each task, while the latter is responsible for ensuring their causality. The modules achieve their objectives through a simple bi-level optimization mechanism with regularization terms. By incorporating MetaCRL into meta-learning, we dynamically eliminate task confounders during the training process. Through extensive evaluations on multiple meta-learning benchmarks, we demonstrate that MetaCRL can significantly improve performance.
Our contributions are as follows: (i) We discover a counterintuitive phenomenon: by increasing the batch size during meta-training, the model’s generalization becomes worse; (ii) We construct an SCM to analyze the phenomenon, finding spurious correlations, named “Task Confounders”, between non-shared features of training tasks and the generic label space of meta-learning; (iii) We propose MetaCRL, a plug-and-play meta-learning causal representation learner, effectively eliminating the task confounders and improving generalization performance; (iv) Extensive empirical analysis showcases the outstanding performance of MetaCRL.
Related Work
@@
Meta-learning aims to construct tasks and learn from them using limited data, thus generalizing to new tasks based on the acquired knowledge. Typical methods can be categorized into two types: optimization-based (Finn, Abbeel, and Levine 2017; Nichol and Schulman 2018; Raghu et al. 2019) and metric-based (Snell, Swersky, and Zemel 2017; Sung et al. 2018; Chen et al. 2020) approaches. They both rely on shared structures to extract meta-knowledge, resulting in remarkable performance on new tasks. However, meta-learning still faces the crisis of performance degradation. Various approaches have been proposed to address this issue, such as adding adaptive noise(Lee et al. 2020), reducing inter-task disparities(Jamal and Qi 2019), limiting the trainable parameters(Yin et al. 2019; Oh et al. 2020), and task augmentation(Yao et al. 2021). These methods overlook the influence between tasks during the training process, which is shown to be crucial in the motivation section. In this study, we focus on the fundamental causes of performance degradation in meta-training tasks and attempt to find a general method to address this problem.
@@
Causal learning is an important branch of machine learning, aiming to understand and infer causal relationships between events. It models the target with a directed acyclic graph and helps in optimizing the model by eliminating confounders in the causal graph. Recent studies (Yang, Zhang, and Cai 2021; Zhang et al. 2020; Nogueira et al. 2022) have already shown that it aids deep learning models in unearthing underlying causal factors. Current research attempts to combine causal knowledge with meta-learning methods to address domain challenges. Yue et al. (Yue et al. 2020) removed performance limitations of pre-trained knowledge through backdoor regulation. Ton et al. (Ton, Sejdinovic, and Fukumizu 2021) utilized causal knowledge to distinguish causes and effects in a bivariate environment with limited data. Jiang et al. (Jiang et al. 2022) used causal graphs to remove undesirable memory effects. While they all combine meta-learning and causal learning, their focus is on addressing problems that differ significantly from ours.
Problem Formulation and Analysis
Notation and Problem Definition
@@
Given a task distribution $p(\mathcal{T})$, the meta-training set $\mathcal{D}_{tr}$ and the meta-testing set $\mathcal{D}_{te}$ are all sampled from $p(\mathcal{T})$ without any class-level overlap. We denote the $N_{tr}$ tasks in a single training batch as $\left\{\tau_{i}\right\}_{i=1}^{N_{tr}}\in\mathcal{D}_{tr}$. Each task $\tau_{i}$ consists of a support set $\mathcal{D}_{i}^{s}=(X_{i}^{s},Y_{i}^{s})=\{(x_{i,j}^{s},y_{i,j}^{s})\}_{j=1}^{N_{i}^{s}}$ and a query set $\mathcal{D}_{i}^{q}=(X_{i}^{q},Y_{i}^{q})=\{(x_{i,j}^{q},y_{i,j}^{q})\}_{j=1}^{N_{i}^{q}}$, where $(x_{i,j}^{\cdot},y_{i,j}^{\cdot})$ represents the sample and the corresponding label, and $N_{i}^{\cdot}$ denotes the number of the samples. Meta-learning utilizes the shared encoder $g$ and the classifier $h$ to learn the above tasks. For simplicity, we denote the meta-learning model as $f_{\theta}=h\circ g$ with the meta-parameter $\theta$ to be learned. Note that the optimal $f_{\theta}$ can serve as general initial priors to help task-specific models adapt quickly to new tasks.
Following (Gordon et al. 2018; Jiang et al. 2022), the objective of meta-learning can be formulated as maximizing the conditional likelihood $\sum\nolimits_{i}{p(Y_{i}^{q}|X_{i}^{q},{\cal D}_{i}^{s},{f_{\theta}})}$. This can be viewed as a bi-level optimization process. In this context, the inner-loop optimizes $p(Y_{i}^{s}|X_{i}^{s},f_{\theta})$ for learning the $i$-th task-specific model $f^{i}_{\theta}$, which can be presented as:
@@
where $\alpha$ is the learning rate. The outer loop optimizes $\sum\nolimits_{i}{p(Y_{i}^{q}|X_{i}^{q},f_{\theta}^{i},{f_{\theta}})}$ for learning $f_{\theta}$, which can be presented as:
@@
where $\beta$ is the learning rate. It is worth noting that $f_{\theta}^{i}$ is obtained by taking the derivative of $f_{\theta}$, so $f_{\theta}^{i}$ can be regarded as a function of $f_{\theta}$. Therefore, $\nabla_{f_{\theta}}\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathcal{L}(Y_{i}^{q},X_{i}^{q},f_{\theta}^{i})$ can be viewed as the second derivative of $f_{\theta}$.
Empirical Evidence
@@
According to Eq. 1 and Eq. 2, we can conclude that meta-learning can be viewed as a multi-task learning process. Meanwhile, a well-generalized meta-learning model encodes the label-related factors for all tasks to obtain rich knowledge. Therefore, intuitively, one might assume that increasing the batch size would enhance the performance. However, our experiments show that this is not always true.
In our empirical evaluation, we begin by sampling a collection of tasks, denoted as $\mathcal{T}_{1}=\left\{\tau_{i}\right\}_{i=1}^{N_{1}}$. Subsequently, we partition the sample sets of all tasks within $\mathcal{T}_{1}$ into two distinct subsets: one allocated for training, denoted as $\mathcal{D}_{1}$, and the other designated for testing, denoted as $\mathcal{D}_{2}$. Following this, we proceed to sample a new task collection $\mathcal{T}_{2}=\left\{\tau_{i}\right\}_{i=1}^{N_{2}}$, ensuring that the tasks do not overlap with $\mathcal{T}_{1}$. We establish MAML as our baseline method and select miniImagenet (Vinyals et al. 2016), Omniglot (Lake, Salakhutdinov, and Tenenbaum 2019), tieredImagenet (Ren et al. 2018) and CIFAR-FS (Bertinetto et al. 2018) as benchmark datasets. During the training phase, we utilize two different batch size settings on $\mathcal{D}_{1}$ to train MAML, i.e., batch size=$B$ and batch size=$2B$, where $B$ represents the commonly used setting on the corresponding dataset, e.g., $B=4$ for miniImagenet and $B=32$ for Omniglot. Subsequently, we evaluate the trained models on $\mathcal{D}_{2}$ and $\mathcal{T}_{2}$, respectively.
Some abnormal outcomes are summarized in Figure 1 (see appendix for full results). Our observations reveal the following: (i) Increasing the number of tasks per batch during training leads to a decrease in test accuracy on $\mathcal{D}_{2}$. It indicates that simply increasing the task quantity does not guarantee an improvement in the model’s performance on the current tasks. Furthermore, this observation suggests that the synergistic enhancement between different tasks is not always evident; in fact, instances of task confounders, where different tasks inhibit each other, may also manifest. (ii) The test accuracy on $\mathcal{T}{2}$ also decreases when the number of tasks per batch during training is increased. In conjunction with observation (i), it becomes evident that the generalization of the meta-learning model is suppressed when there are task confounders among multiple tasks.
Causal Analysis and Motivation

### v4

@@ Abstract
Meta-learning enables rapid generalization to new tasks by learning knowledge from various tasks. It is intuitively assumed that as the training progresses, a model will acquire richer knowledge, leading to better generalization performance. However, our experiments reveal an unexpected result: there is negative knowledge transfer between tasks, affecting generalization performance. To explain this phenomenon, we conduct Structural Causal Models (SCMs) for causal analysis. Our investigation uncovers the presence of spurious correlations between task-specific causal factors and labels in meta-learning. Furthermore, the confounding factors differ across different batches. We refer to these confounding factors as “Task Confounders”. Based on these findings, we propose a plug-and-play Meta-learning Causal Representation Learner (MetaCRL) to eliminate task confounders. It encodes decoupled generating factors from multiple tasks and utilizes an invariant-based bi-level optimization mechanism to ensure their causality for meta-learning. Extensive experiments on various benchmark datasets demonstrate that our work achieves state-of-the-art (SOTA) performance. The code is provided in https://github.com/WangJingyao07/MetaCRL.
1 Introduction
Meta-learning aims to develop models that can be rapidly transferred to previously unseen tasks. To achieve this, it first learns from diverse tasks to obtain models with high learning capacities. Then, it fine-tunes these models with little data from unseen tasks to obtain the desired ones. Recently, meta-learning has been widely applied in various fields, e.g., affective computing  Li et al. (2023), image classification  Qiang et al. (2023), and robotics  Schrum et al. (2022).
During the training phase, each batch consists of a series of randomly sampled $N$-way $K$-shot tasks, where $N$ denotes the number of classes per task and $K$ denotes the number of samples per class. The samples in each task are divided into a support set and a query set. Then, meta-learning models are trained in a bi-level optimization manner  Wang et al. (2021, 2023). In brief, at the first level, the desired model for each task is fine-tuned by training on the support set using the meta-learning model. At the second level, the meta-learning model is learned using the query sets from all training tasks and the corresponding expected models for each task. Therefore, a widely adopted hypothesis is that as training progresses, the meta-learning model will acquire richer knowledge that can be transferred well to downstream tasks, achieving better performance  Rivolli et al. (2022).
Figure 1: Knowledge transfer to a specific test task. For both positive knowledge transfer ($\mathcal{R}_{i,j}<1$) and negative knowledge transfer ($\mathcal{R}_{i,j}>1$), an exemplar task is shown. Here, we simply use the $\mathcal{R}_{i,j}$ threshold to classify the knowledge transfer as positive or negative. See Subsection 3.2 and Appendix F for more details.
@@ 1 Introduction
(b) meta-learning process
@@
Figure 2: Structural Causal Models (SCM) regarding two tasks $\tau_{i}$ and $\tau_{j}$, where $(X_{i},Y_{i})$ and $(X_{j},Y_{j})$ are the samples and corresponding labels of these tasks. The solid line means the true causal correlation, and the dotted line means the spurious correlation. (a) is constructed based on the ground-truth causal mechanism, while (b) can be viewed as the inverse process of the generating mechanism.
However, our toy experiments reveal a conflicting phenomenon, i.e., the knowledge learned from the training tasks may be harmful to the unseen test tasks (See Subsection 3.2 for more details). Specifically, we first randomly sample 400 tasks from miniImageNet dataset Vinyals et al. (2016) and divide them into a training set and a test set. Then, we define a metric $\mathcal{R}_{i,j}$ to evaluate whether the meta-learning model trained on the training tasks can perform better on the test task, i.e., quantify the knowledge transfer performance from the training tasks to each test task. If $\mathcal{R}_{i,j}<1$, the learned knowledge from the training task can help improve the model performance on the test task (positive knowledge transfer), while $\mathcal{R}_{i,j}>1$ implies the learned knowledge is harmful to the test task (negative knowledge transfer). We use MAML Finn et al. (2017) as the baseline and record the score of $\mathcal{R}_{i,j}$ in the middle of training  Fifty et al. (2020); Abdollahzadeh et al. (2021). Figure 1 shows the results. Ideally, all the knowledge transfer between tasks should be positive, i.e., $\mathcal{R}_{i,j}<1$. The results show that there always exists negative knowledge transfer between tasks.
To explore the reasons behind this phenomenon, we propose using causal theory for analysis (See Subsection 3.3 for details). We begin by constructing Structural Causal Models (SCMs) for the training phase of ML, as shown in Figure 2. In the SCMs, ${\rm{A}}^{i}$ and ${\rm{A}}^{j}$ are the distinct causal factors of task $\tau_{i}$ and task $\tau_{j}$, and ${\rm{B}}^{i,j}$ means the shared causal factors of these two tasks. Meanwhile, causal factors can be considered as different semantics of the data, e.g., color and shape, also considered as generating factors used for data generation Zimmermann et al. (2021). Since meta-learning performs joint learning on all the training tasks, it acquires all the causal factors. Thus, the non-overlapping causal factors ${\rm{A}}^{i}$ of $\tau_{i}$ may cause spurious correlations with $\tau_{j}$, and ${\rm{A}}^{j}$ holds the same with $\tau_{i}$. These misleading correlations between training tasks will introduce bias into the learned knowledge and ultimately affect generalization, which is called “task confounder”.
@@
To address this issue, we propose a plug-and-play meta-learning causal representation learner (MetaCRL) to encode decoupled causal knowledge, thereby eliminating task confounders. It consists of two modules: the disentangling module and the causal module. The former aims to extract generating factors across all tasks and provide a subset of factors relevant to each task, while the latter is responsible for ensuring their causality. The modules achieve their objectives through a simple bi-level optimization mechanism with regularization terms. By incorporating MetaCRL into meta-learning, we dynamically eliminate task confounders during the meta-training process. Through extensive evaluations of multiple meta-learning benchmarks, we demonstrate that MetaCRL can significantly improve performance.
In summary, our contributions are as follows:
  • We discover a counterintuitive phenomenon: there is negative knowledge transfer between tasks, resulting in reduced model generalization performance.
  • We construct an SCM to analyze the phenomenon with causal theory, finding spurious correlations, named “Task Confounders”, between non-shared causal factors of the meta-training tasks and the label space.
  • We propose MetaCRL, a plug-and-play meta-learning causal representation learner to eliminate task confounders, thus improving generalization performance.
  • Extensive experiments on various scenarios demonstrate the outstanding performance of our MetaCRL.
2 Related Work
Meta-learning
@@ 2 Related Work
aims to learn general knowledge from various training tasks, and then generalize to new tasks based on the acquired knowledge. Typical methods can be categorized into two types: optimization-based  Finn et al. (2017); Nichol and Schulman (2018); Guo et al. (2024) and metric-based  Snell et al. (2017); Sung et al. (2018); Chen et al. (2020) methods. They both rely on shared structures and bi-level learning mechanisms to learn general knowledge, resulting in remarkable performance on new tasks. However, meta-learning still faces the crisis of performance degradation. Various approaches have been proposed to address this issue, such as adding adaptive noise Lee et al. (2020), reducing inter-task disparities Jamal and Qi (2019), limiting the trainable parameters Yin et al. (2019); Oh et al. (2020), and task augmentation Yao et al. (2021). Despite alleviating performance degradation, they ignore the interaction between tasks, which is shown to be crucial in Section 3. In this study, we analyze the knowledge transfer effects between different training tasks with causal theory, and focus on the fundamental causes of performance degradation in meta-learning.
Causal learning
@@
aims to explore the causal relationships between variables in machine learning, modeling the target with a directed acyclic graph, also known as a causal model. It has been shown to aid models in unearthing underlying causal factors  Yang et al. (2021); Zhang et al. (2020); Nogueira et al. (2022). Current research attempts to combine causal knowledge with meta-learning methods to address domain challenges. Yue et al.  Yue et al. (2020) removed performance limitations of pre-trained knowledge through backdoor regulation. Ton et al.  Ton et al. (2021) utilized causal knowledge to distinguish causes and effects in a bivariate environment with limited data. Jiang et al.  Jiang et al. (2022) used causal graphs to remove undesirable memory effects. While they all combine meta-learning and causal learning, their focus is on addressing problems that differ from ours.
3 Problem Formulation and Analysis
In this section, we first present the notation and problem definition of meta-learning. Next, we conduct experiments to evaluate the interaction between different tasks and illustrate the empirical evidence, i.e., the knowledge learned from the training tasks may be harmful to the unseen test tasks, reducing generalization performance. Finally, we construct SCMs to explore the reasons behind the empirical evidence.
3.1 Preliminaries
@@
Given a task distribution $p(\mathcal{T})$, the meta-training dataset $\mathcal{D}_{tr}$ and the meta-test dataset $\mathcal{D}_{te}$ are all sampled from $p(\mathcal{T})$ without class-level overlap. During the training phase of ML, each batch contains $N_{tr}$ tasks, denoted as $\left\{\tau_{i}\right\}_{i=1}^{N_{tr}}\in\mathcal{D}_{tr}$, and each task $\tau_{i}$ consists of a support set $\mathcal{D}_{i}^{s}=(X_{i}^{s},Y_{i}^{s})=\{(x_{i,j}^{s},y_{i,j}^{s})\}_{j=1}^{N_{i}^{s}}$ and a query set $\mathcal{D}_{i}^{q}=(X_{i}^{q},Y_{i}^{q})=\{(x_{i,j}^{q},y_{i,j}^{q})\}_{j=1}^{N_{i}^{q}}$, where $(x_{i,j}^{\cdot},y_{i,j}^{\cdot})$ represents the sample and the corresponding label, and $N_{i}^{\cdot}$ denotes the number of the samples. The meta-learning model $f_{\theta}=h\circ g$ utilizes the feature encoder $g$ and the classifier $h$ to learn the above tasks.
The learning mechanism of meta-learning is regarded as a bi-level optimization process. At the first level, it fine-tune the desired model $f_{\theta}^{i}$ for task $\tau_{i}$ by training on the support set $\mathcal{D}_{i}^{s}$ using the meta-learning model $f_{\theta}$, presented as:
@@
where $\alpha$ is the learning rate. At the second level, the meta-learning model $f_{\theta}$ is learned using the query sets $\mathcal{D}^{q}$ from all training tasks and the expected models for each task:
@@
where $\beta$ is the learning rate. Note that $f_{\theta}^{i}$ is obtained by taking the derivative of $f_{\theta}$, so $f_{\theta}^{i}$ can be regarded as a function of $f_{\theta}$. Therefore, the update of $f_{\theta}$ mentioned in Eq.2 can be viewed as calculating the second derivative of $f_{\theta}$.
3.2 Empirical Evidence
@@
From above and Wang et al. (2021), meta-training on one batch can be viewed as a multi-task learning process. Meanwhile, a well-learned model should contain knowledge of all training tasks. Therefore, intuitively, one might assume that as training progresses, the meta-learning model will acquire richer knowledge (related to all tasks) and transfer better to downstream tasks, achieving great generalization. However, our toy experiments reveal that this is not always true.
Before introducing the toy experiments, we first present a method to quantify the influence of transferring knowledge learned from one task to the target task. For task $\tau_{i}$, the model $f_{\theta}$ uses the support set $\mathcal{D}_{i}^{s}$ to obtain $f_{\theta}^{i}$ via Eq.1. Here, $f_{\theta}^{i}$ is considered to integrate the knowledge of task $\tau_{i}$ into $f_{\theta}$. Then, for task $\tau_{j}$, we first obtain the model $f_{\theta}^{j,1}$ by training $f_{\theta}^{i}$ on the support set $\mathcal{D}_{j}^{s}$, and then obtain the model $f_{\theta}^{j,2}$ by training $f_{\theta}$ on $\mathcal{D}_{j}^{s}$. Next, we calculate their losses on the query set $\mathcal{D}_{j}^{q}$, expressed as $\mathcal{L}(\mathcal{D}_{j}^{q},f_{\theta}^{j,1})$ and $\mathcal{L}(\mathcal{D}_{j}^{q},f_{\theta}^{j,2})$, respectively. Finally, we calculate the ratio between these two losses, denoted as $\mathcal{R}_{i,j}$, which quantifies the performance of knowledge transfer from task $\tau_{i}$ to task $\tau_{j}$. Thus, we have:
$\begin{array}[]{l}\mathcal{R}_{i,j}=\frac{\mathcal{L}(\mathcal{D}_{j}^{q},f_{\theta}^{j,1})}{\mathcal{L}(\mathcal{D}_{j}^{q},f_{\theta}^{j,2})}\end{array}$ (3)
if $\mathcal{R}_{i,j}<1$, it means that task $\tau_{i}$ has a positive knowledge transfer effect on task $\tau_{j}$. On the other hand, if $\mathcal{R}_{i,j}>1$, it indicates the negative knowledge transfer effect of $\tau_{i}$ on $\tau_{j}$.
Next, we conduct experiments based on the quantitative method described above. We first randomly sample 400 tasks from miniImageNet dataset, which are divided into a training set of 300 tasks and a test set of 100 tasks. Then, we use MAML as the baseline to calculate the score of $\mathcal{R}_{i,j}$ from the training tasks to each test task in the middle of training.
Figure 1 shows the histograms of the knowledge transfer in the training phase of meta-learning along with exemplar tasks. From the results, we observe that as training proceeds, although the knowledge transfer effects become more and more positive, there always exists negative knowledge transfer between different tasks. It indicates that the training process of meta-learning cannot always obtain effective knowledge for unseen test tasks, and the aforementioned intuitive hypothesis is limited. Note that we also conduct experiments under various different settings, including using multiple meta-learning baselines, using different datasets, and training on multiple tasks simultaneously (the effect of multiple training tasks to a single test task), the impact of negative knowledge transfer always exists. More details and the full results are provided in Appendix F.
3.3 Causal Analysis and Motivation